A Bibliography of Numerical Analysis Publications, 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 December 2023 Version 1.70

Title word cross-reference

0 [Bor14].

1 [SS18a].

2 [SS18b]. 2010 [BJ12]. 2011 [ABIL13].

3 [DW11, Kiu13].

5th [Gre16b].

60th [BBB^{+}13].

8 [HLR14].

978 [Bor14]. 978-2-9700636-5-0 [Bor14]. accelerators [KBD10]. Algebra [AI14, BR11, BV18, BCS14, CK14, GH17, Gru14, Hog14, Lay14, MM18, Sin13, Dat10, Dav12, KG11]. Algorithm [Wat11]. algorithmic [Nau11]. Algorithmische [KW12]. Algorithms [BCS14, Mul16, BM17, KG11, Mag10, SS18a]. Analysis [ABIL13, BR11, Blo14, BJ12, CF13, GMQS17, GHLS12, Ham12, HJ12, LR12, Mil14, NF14, RZ16, SS14, AAKS10, BMPW10, Bra14, BF10, FM10, GS12, How17, Lan10, Lau12, MMS17, Moi10, Oña10, Sco11, BC10]. analytical [BBB^{+}13, GS12]. APIs [Meh15]. Applicable [GMQS17]. Applications [ABIL13, BCS14, GM10, HPS13, Lay14, Mol12, Sch17, B´ 10, BLW11, Che11, Dat10, Dav12, FM10, KM10a, KM11, LB13, RZ15, Riz12, SS18b]. Applied [BV18, Ham12, Reb12]. approach [KG12, Sha12]. Approximate [LR12]. Approximation [GMR11, GMQS17, HH13, Tre13, Coh11, GV11]. Approximations [ABIL13, AH12]. aproximacão [GV11]. Archive [Bor14]. Arithmetic [KM10b, BM17]. art [Nau11]. assisted [NPW19]. Asymptotic [Tem15].

B [HH13]. B-splines [HH13]. Backward [CF13]. Band [KW12]. Basic [Zha11]. beginners [HLR14]. Berlin [Gre16b]. big [Kut13]. bioengineering [KM10a, KM11]. birth [BC10]. Birthday [BBB^{+}13]. Book [Bor14]. Borwein [BBB^{+}13]. Boundary [SH11].

C [EK12, Sha12]. calculation [BMPW10]. Carlo [DP10, GT13, RC10]. Cetraro [ABIL13]. Chasing [Wul12]. Chebyshev [Boy14]. Classical [AAKS10, SS18b]. Claude [Bor14]. coding [SS18a]. Collocation [Sch17]. combinatorial [Mol12]. Commentaries [BS13, BS14a, BS14b, KO10]. companion [Den13, Gre16a]. complete [Meh15]. Complex [GMQS17, Kut13]. Comprehensive [AWH13]. Computation [Bee17, GMR11, RZ16, Kut13]. Computational [BBB^{+}13, HRSW13, How17, Lau12, ŠH12, War13, Bra14, Nah11, Tra10, VF10]. Computations [GV13, Wat10, Tuc11]. Computer [BM17, NPW19, BS10, Dav12, Nau11, SS18a]. Computer-assisted [NPW19]. Computers [Kne16, Kne17]. Computing [GW12, NF14, BS11, EK12, Gus11, HH14, KBD10, Mag10, Meh15, OR10, Poz13, QSG10, QSG14, Ros13, Ros14, TLNC10, VF10]. Conference [Gre16b, RZ15]. Congress [FVJT10]. Consumer [Wal14]. Continuous [KD14]. Control [KD14]. cookbook [Ros14]. Coq [BM17]. Cosmology [Gre16a]. Course [GH17, AG11, Reb12]. courses [FM10]. crunching [Nah11].

Data [Bra14, HPS13, Kut13, MMS17, Riz12, Ros13, Ros14]. Data-driven [Kut13]. Deconvolution [LR12]. Derivative [HRSW13]. destin [Bor14, PR12]. diferenciais [GV11]. difference [LQT18]. Differential [AI14, NPW19, Sch17, Dav11a, GV11, GH10, GS12, LQT18, SS18a, SS18b, SCM12]. differentiating [Nau11]. differentiation [Nau11]. Digital [DP10]. dimensional [ÖM13]. discrepancy [DP10]. douloureux [Bor14, PR12]. driven [Kut13]. Durham [BJ12]. dynamic [KA11]. Dynamical [CK14].

Ed [Bor14]. Eigenvalue [Saa11].

Eigenvalues [CCAL12]. Einstein [BS10]. Element [HRSW13, Che11, Dav11a, GV11, LB13, LQT18, Oña10]. Elementary [Mul16]. elementos [GV11]. Elements [TLNC10, BWS11, Joh11, ÖM13]. Engineering [Blo14, B´ 10, Che11, Chr10, Kiu10a, Kiu10b, Kiu13, Moi10, VF10]. Engineers [Mil14, Ant12, Bra14]. equacões [GV11]. Equations [ABIL13, NPW19, Sch17, BS10, Boy14, Dav11a, GV11, GH10, GS12, LQT18, SS18a, SS18b, SCM12]. Error [CF13, BWS11]. estimation [BWS11]. État [Bor14]. Euler [GW12]. event [Riz12]. exercises [Gau16]. expansions [Chr10]. experienced [HLR14]. Exploratory [MMS17].

FE [ÖM13]. fiction [Nah11]. figures

[TLNC10]. Finance

[GMS11, HRSW13, B´ 10]. Finite

[BWS11, HRSW13, KM10b, Che11, Dav11a, GV11, LB13, LQT18, ÖM13, Oña10]. finite-element [GV11]. finitos [GV11]. first [AG11, Reb12]. Fitting [HPS13]. floating [BM17]. floating-point [BM17]. formal [BM17]. Fortran [HH14]. Foundation [RZ15]. Foundations [GT13]. Four [FM10]. Fourier [Has18]. fractional [SS18a, SS18b]. frames [FM10]. Francis [Wat11]. frequency [FM10]. Frontiers [BJ12]. Function [Bee17]. Functions [Chr10, Mul16, SH11]. Fundamentals [Gus11, Mag10, Moi10, Tra10, Wat10, Che11].

G. [KO10]. Galerkin [GW12]. Gautschi [BS13, BS14a, BS14b]. génie [PR12, Bor14]. Geometry [AI14, Fie11]. German [DW11]. Germany [Gre16b]. Gradimir [GMR11]. Graduate [CF13]. Graph [KG11]. Graphs [Fie11, Mol12]. Green [SH11]. grid [Mag10]. Guide [Wal14, AWH13, HH17, HLR14, Meh15]. Guided [SV15].

Haifa [RZ15]. Hamilton [ABIL13]. Handbook [Bee17, Hog14, Sch15, Ste11]. hands [Ros14]. hands-on [Ros14]. harmonic [FM10]. Harmonics [AH12]. heavens [Wul12]. Heisenberg [Raz11]. high [Ros13, Ros14]. high-performance [Ros13, Ros14]. Honor [GMR11, BBB^{+}13].

ICMS [FVJT10, Gre16b]. II [AI14]. image [FM10]. Imaging [Sch15]. Implementation [Mul16, LB13]. Infinite [RZ15]. initial [GH10]. Insight [VF10]. integer [SS18b]. Integrals [Tem15]. integration [DP10, SS18a, SS18b]. Intelligent [AI14]. interactive [Ros13, Ros14]. International [FVJT10, Gre16b]. Introducão [GV11]. Introducing [RC10]. Introduction [AH12, BV18, CF13, Ett10, Ham12, JMR14, NF14, Att12, BWS11, BS11, Dav11a, KG12, LQT18, Nau11, ÖM13, Tuc11, VF10, War13, GV11, SS18a]. IPython [Ros13, Ros14]. ISBN [Bor14]. Israel [RZ15]. Italy [ABIL13].

Jacobi [ABIL13]. Japan [FVJT10]. Java [Alt12]. Jean [Bor14]. Jean-Claude [Bor14]. Joint [Riz12]. Jonathan [BBB^{+}13]. July [Gre16b].

Kobe [FVJT10].

language [KG11]. Laplacian [Mol12]. Large [Saa11]. Learn [Ros13]. Learning [Hil15, Ros13]. Least [BV18, HPS13]. Lecture [SS14]. Library [Bee17]. Linear [AI14, BR11, BV18, BCS14, CK14, Dav12, GH17, Gru14, Hog14, Lay14, MM18, Sin13, Dat10, KG11, Oña10]. longitudinal [Riz12].

macromixing [BMPW10]. Maple [GS12]. master [Meh15]. Mastering [Meh15]. MathCW [Bee17]. Mathematica [SS14]. Mathematical [AWH13, Bee17, FVJT10, GT13, RZ16, Sch15, SV15, BS11, Chr10, Nah11, RZ15, FVJT10, Gre16b]. Mathematical-Function [Bee17]. Mathematics [RZ16, BBB^{+}13, DW11, Reb12]. Mathematik [DW11, Her11]. MATLAB [Alt12, Att12, B´ 10, Dav11b, Ett10, Gau16, Gre16a, GS12, HH17, HLR14, Joh11, KG12, KM10a, KM11, MMS17, VF10, War13, QSG10, QSG14, SS15]. MATLAB(R) [Kiu10a, KA11]. Matrices [BV18, Fie11, GM10, CCAL12, Mol12]. Matrix [BR11, GV13, Gru14, HJ12, Wat10, Zha11, Lau12, Mol12]. May [RZ15]. measure [Wul12]. mechanics [Raz11]. media [Gre16a]. Memory [GMQS17]. method [BWS11, Che11, Dav11a, GV11, LB13, ÖM13, Oña10]. Methoden [KW12]. Methods [Wal14, AWH13, CF13, GMS11, GT13, HRSW13, KD14, NPW19, Saa11, Sch15, Sch17, SV15, Ste11, Tem15, AAKS10, Ant12, AG11, Bra14, Coh11, FM10, GH10, GS12, How17, KG12, KM10a, KM11, Kiu10a, Kiu10b, Kiu13, Kut13, LQT18, Mon11, RC10, Sha12, ŠH12]. método [GV11]. Micro [BMPW10]. Micro- [BMPW10]. Milovanović [GMR11]. Modeling [SS15, HH13, Kut13]. Models [Gru14, LR12, Riz12]. Modern [GW12, AAKS10, BS11, HH14]. Moments [GM10]. Monte [DP10, GT13, RC10]. multicore [KBD10]. Multiscale [GHLS12].

nets [DP10]. neuroscience [Tra10]. Notes [SS14]. Number [KM10b, Nah11]. Number-crunching [Nah11]. Numbers [Kne16, Kne17]. numérica [GV11]. Numerical [ABIL13, Wal14, Ant12, BS10, Blo14, BJ12, BF10, Coh11, CF13, Dat10, GMS11, GHLS12, GH10, Ham12, HH14, KM10a, KM11, Kiu10a, Kiu10b, Kiu13, KD14, Lan10, LR12, LQT18, Mil14, Mon11, NPW19, NF14, Saa11, SS18a, SS18b, Sco11, Ste11, SS14, AAKS10, AG11, BMPW10, Boy14, BC10, GV11, GS12, How17, KG12, Moi10, Ros13, Ros14, Sha12, DW11]. numerics [Tuc11]. Numerische [Her11, DW11].

o [GV11]. object [Sha12]. object-oriented [Sha12]. Octave [QSG10, QSG14]. One [ÖM13]. One-dimensional [ÖM13]. operators [RZ15]. Optimization [Wal14, GMS11]. oracles [Boy14]. ordinary [GH10]. oriented [Sha12]. Orthogonal [Gau16]. other [Boy14].

pack [Gre16a]. parciais [GV11]. Partial [NPW19, Sch17, Dav11a, GV11, GS12, SS18a, SS18b]. PDEs [SS18b, Sha12]. Performance [BLW11, Ros13, Ros14]. perturbation [Boy14]. Phenomenology [LR12]. photonics [War13]. Physical [SV15]. physicien [Bor14, PR12]. Physicists [AWH13, ŠH12]. physics [B´ 10, Chr10, Nah11]. point [BM17]. polynomial [Boy14]. polynomials [Gau16]. Pont [Bor14]. Portable [Bee17]. Portuguese [GV11]. practical [Att12]. Practice [Has18, Tre13, AAKS10].

Precision [KM10b]. Pricing [HRSW13]. primer [Dav11b, Lan12]. probability [Dav12]. problem [Att12]. Problems [GHLS12, KD14, Saa11, SH11, GH10, Nah11]. Proceedings [FVJT10, Gre16b, RZ15]. products [RZ15]. programmers [Meh15]. Programming [Bee17, Alt12, Att12, Hil15, JMR14, Lan12, Ros13]. programs [Nau11]. Progress [GMQS17]. Proofs [NPW19, BM17]. proxy [Boy14]. Python [Ste14, Hil15, Kiu10b, Kiu13, Lan12, Meh15, Ros13, Ros14].

Q. [GMQS17]. Quadrature [GM10]. Quantification [SS15]. Quantitative [HRSW13]. quantum [Raz11]. quasi [DP10]. quasi-Monte [DP10].

R [Blo14, Den13, EK12, How17, JMR14, Riz12, RC10, SS18a, Sch17, SCM12]. race [Wul12]. Rahman [GMQS17]. recipes [Ros14]. relativity [BS10]. research [RZ15]. Results [Zha11]. Review [Bor14]. rigorous [Tuc11]. Ritz [Bor14, GW12, PR12]. rootfinders [Boy14]. Routines [AI14].

Sage [AI14]. Science

[Blo14, RZ15, Dav12, Nah11, Ros14, VF10]. Sciences [SV15]. Scientific [KBD10, NF14, QSG10, QSG14, BLW11, Gus11, Hil15, JMR14, Kut13, Lan12, Meh15, OR10, Poz13, TLNC10]. Scientists [Mil14, Ant12, Bra14, Ste14]. Second [GH17]. secrets [Alt12]. Selected [BS13, BS14a, BS14b, KO10]. September [FVJT10]. sequences [DP10]. series [Boy14]. sharpen [Ros14]. short [FM10, Tuc11]. signal [FM10]. Simulation [GT13, KA11, BMPW10, JMR14]. Simulink [HLR14]. Simulink(R) [KA11]. Sinc [Ste11]. Sion [Bor14]. skills [Ros14]. Software [Bee17, FVJT10, Gre16b]. solution [LQT18]. solutions [Gau16]. Solving [AI14, Boy14, Sha12, SCM12, Att12, BS10]. space [SS18a, SS18b]. spaces [Chr10]. Sparse [Has18]. Sphere [AH12]. Spherical [AH12]. Spline [Sch17]. splines [HH13]. Squares [BV18, HPS13]. Statistical [EK12, Bra14, KM10a, KM11]. statisticians [Lan10]. Statistics [BR11, Mon11, Oña10]. Step [Sin13]. Stewart [KO10]. Stochastic [GT13, KD14, SS15]. Structural [Oña10]. student [Den13]. style [Joh11]. system [BM17]. Systems [CK14, KM10b, KA11, Kut13].

tables [TLNC10]. taming [Nah11]. Techniques [BCS14, Zha11]. technologies [Mag10]. their [RZ15]. théoricien [Bor14, PR12]. Theory [GMQS17, Has18, Tre13, Zha11, AAKS10, LB13, Mag10, Mol12]. Third [FVJT10]. Time [KD14, FM10, Riz12]. time-frequency [FM10]. time-to-event [Riz12]. tools [Chr10, Meh15]. Tour [SV15]. transcendental [Boy14]. Transform [Has18]. Traveling [GS12]. tuning [BLW11]. Turbulence [LR12].

Uncertainty [SS15]. Undocumented [Alt12]. unified [Sha12]. Unit [AH12]. unruly [Nah11]. updated [HLR14]. users [HLR14]. Using [Bee17, Blo14, JMR14, Meh15].

V [GMR11]. Valais [Bor14]. Validated [Tuc11]. validation [OR10]. Vallesia [Bor14]. Value [SH11, GH10]. Vectors [BV18]. Venus [Wul12]. Verification [NPW19, OR10]. verifying [BM17]. via [GV11]. Viewpoint [CF13]. visualization [Ros13, Ros14]. Volume [BS13, BS14a, BS14b, SS18a, SS18b].

W [KO10]. Walter [BS13, BS14a, BS14b]. Walther [Bor14, PR12]. wave [GS12]. wavelets [FM10]. Works [BS13, BS14a, BS14b, KO10]. workshop [RZ15].

XML [Poz13].

References

Ackleh:2010:CMN

[AAKS10] Azmy S. Ackleh, Edward James Allen, Ralph Baker Kearfott, and Padmanabhan Seshaiyer. Classical and modern numerical analysis: theory, methods and practice. Chapman and Hall/CRC numerical analysis and scientific computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4200-9157-3 (hardcover). xix + 608 pp. LCCN QA297 .C53 2010. Theory, methods and practice.

Achdou:2013:HJE

[ABIL13] Yves Achdou, Guy Barles, Hitoshi Ishii, and Grigory L. Litvinov, editors. Hamilton–Jacobi Equations: Approximations, Numerical Analysis and Applications: Cetraro, Italy 2011, volume 2074 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LN-MAA2. ISBN 3-642-36432-2 (print), 3-642-36433-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2074; QA3 .L28 no. 2074; QA316 .C56 2011.

                                              URL http://link.springer.com/
                                              book/10.1007/978-3-642-36433-4;
                                               http://www.springerlink.com/content/
                                              978-3-642-36433-4.

Ascher:2011:FCN

[AG11] Uri M. (Uri M.) Ascher and Chen Greif. A first course in numerical methods, volume 7 of Computational science and engineering. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2011. ISBN 0-89871-997-6. xxii + 552 pp. LCCN QA297 .A748 2011.

URL http://www.loc.gov/catdir/
enhancements/fy1111/2011007041-
b.html;   http://www.loc.gov/catdir/
enhancements/fy1111/2011007041-
d.html;   http://www.loc.gov/catdir/
enhancements/fy1111/2011007041-
t.html.

Atkinson:2012:SHA

[AH12] Kendall Atkinson and Weimin Han.

Spherical Harmonics and Approximations on the Unit Sphere: an Introduction, volume 2044 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-25982-0 (print), 3-642-25983-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 244 pp. LCCN QA3 .L28 no. 2044; QA406 .A85 2012.

URL http://link.springer.com/
book/10.1007/978-3-642-25983-8;
 http://www.springerlink.com/content/
978-3-642-25983-8;   http://www.
springerlink.com/content/u58550t8417n/

.

Anastassiou:2014:IRI

[AI14] George A. Anastassiou and Iuliana F.

Iatan. Intelligent Routines II: Solving Linear Algebra and Differential Geometry with Sage, volume 58 of Intelligent systems reference library. Springer International Publishing, Cham, Switzerland, 2014. ISBN 3-319-01966-X, 3-319-01967-8 (e-book). ISSN 1868-4394. xiv + 306 pp. LCCN QA614 .A63 2014. URL http://d-

                                        nb.info/1045077038/34;   http:
                                        //nbn-resolving.de/urn:nbn:de:
                                        1111-201312062402;   http://www.
                                        springerlink.com/content/978-3-
                                        319-01967-3.

Altman:2012:USM

[Alt12] Yair M. Altman. Undocumented secrets of MATLAB–Java programming. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-6904-9 (electronic bk.), 1-4398-6903-0 (hardback), 1-4398-6903-0. xxi + 663 + 16 pp. LCCN QA297 .A544 2012.

Antia:2012:NMS

[Ant12] H. M. Antia. Numerical methods for scientists and engineers, volume 2 of Texts and Readings in Physical Sciences. Hindustan Book Agency, New Delhi, New Delhi, third edition, 2012. ISBN 93-80250-40-1 (hardcover). xxxii + 855 pp. LCCN TA335 .A58 2012.

Attaway:2012:MPI

[Att12] Stormy Attaway. MATLAB: a practical introduction to programming and problem solving. Butterworth-Heinemann, Waltham, MA, USA, second edition, 2012. ISBN 0-12-385081-9. xx + 518 pp. LCCN QA297 .A87 2012.

Arfken:2013:MMP

[AWH13] George B. (George Brown) Arfken, Hans-Jürgen Weber, and Frank E. Harris. Mathematical Methods for Physicists: a Comprehensive Guide.

Elsevier Academic Press, Amsterdam, The Netherlands, seventh edition, 2013. ISBN 0-12-384654-4 (hardcover). xiii + 1205 pp. LCCN QA37.3 .A74 2013.

BaezLopez:2010:MAE

[B´ 10] David Báez López. MATLAB with applications to engineering, physics and finance. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-0697-7 (hardcover). xiv + 412 pp. LCCN QA297 .B28 2010.

Bailey:2013:CAM

[BBB^{+}13] David H. Bailey, Heinz H.

Bauschke, Peter Borwein, Frank Garvan, Michel Théra, Jon D. Vanderwerff, and Henry Wolkowicz, editors. Computational and analytical mathematics: in honor of Jonathan Borwein’s 60th Birthday, volume 50 of Springer proceedings in mathematics and statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-7620-8, 1-4614-7621-6 (e-book). ISSN 2194-1009. xv + 701 pp. LCCN QA241.

URL  http://public.eblib.com/
choice/publicfullrecord.aspx?p=
1466708;   http://swb.eblib.com/
patron/FullRecord.aspx?p=1466708;
 http://www.myilibrary.com?id=
547562.

Bultheel:2010:BNA

[BC10] Adhemar Bultheel and Ronald Cools, editors. The birth of numerical analysis. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2010. ISBN 981-283-625-X. LCCN QA297 .B54 2010.

Bronson:2014:LAA

[BCS14] Richard Bronson, Gabriel B. Costa, and John T. Saccoman. Linear Algebra: Algorithms, Applications, and Techniques. Elsevier Academic Press, Amsterdam, The Netherlands, third edition, 2014. ISBN 0-12-391420-5 (paperback), 0-12-397811-4 (e-book). xi + 519 pp. LCCN QA184.2 .B76 2014.

                                        URL http://www.sciencedirect.
                                        com/science/book/9780123914200.

Beebe:2017:MFC

[Bee17] Nelson H. F. Beebe. The Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-64109-3 (hardcover), 3-319-64110-7 (e-book). xxxvi + 1114 pp. LCCN QA75.5-76.95.

                                        URL http://www.springer.com/us/
                                        book/9783319641096.

Burden:2010:NA

[BF10] Richard L. Burden and J. Douglas Faires. Numerical analysis. Cengage Learning, Boston, MA, USA, nineth edition, 2010. ISBN 0-538-73351-9. ???? pp. LCCN ????

Blowey:2012:FNA

[BJ12] James Blowey and Max Jensen, editors. Frontiers in Numerical Analysis — Durham 2010, volume 85 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSA6. ISBN 3-642-23913-7 (print), 3-642-23914-5 (ebook). ISSN 1439-7358. LCCN ????

URL http://link.springer.com/
book/10.1007/978-3-642-23914-4;
http://www.springerlink.com/content/
978-3-642-23914-4. Proceedings of

the Twelfth LMS–EPSRC Summer School in Computational Mathematics and Scientific Computation held at the University of Durham, UK, 25–31 July 2010.

Bloomfield:2014:URN

[Blo14] Victor A. Bloomfield. Using R for Numerical Analysis in Science and Engineering. Chapman and Hall/CRC the R series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2014. ISBN 1-4398-8448-X (hardcover). xxii + 335 pp. LCCN Q183.9 .B56 2014.

URL http://www.crcnetbase.com/
isbn/9781439884492.

Bailey:2011:PTS

[BLW11] David H. Bailey, Robert F. Lucas, and Samuel Watkins Williams, editors. Performance tuning of scientific applications, volume 11 of Chapman and Hall/CRC computational science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-1569-0 (hardback). ???? pp. LCCN Q183.9 .P47 2011.

Boldo:2017:CAF

[BM17] Sylvie Boldo and Guillaume Melquiond.

Computer arithmetic and formal proofs: verifying floating-point algorithms with the Coq system. ISTE Press, London, UK, 2017. ISBN 1-78548-112-6, 0-08-101170-9 (e-book). LCCN QA76.9.C62. URL http://

iste.co.uk/book.php?id=1238.

Bockhorn:2010:MMA

[BMPW10] Henning Bockhorn, Dieter Mewes, Wolfgang Peukert, and Hans-Joachim Warnecke, editors. Micro- and macromixing: analysis, simulation and numerical calculation. Heat and mass transfer. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-04549-9, 3-642-04548-0. xi + 346 pp. LCCN TP156.M5 M537 2010.

Borrelli:2014:BRB

[Bor14] Arianna Borrelli. Book review: Le destin douloureux de Walther Ritz (1878–1909), physicien théoricien de génie, Jean-Claude Pont (Ed.). Vallesia, Archive de l’État du Valais, Sion (2012), ISBN 978-2-9700636-5-

0. Historia Mathematica, 41(1):107– 110, February 2014. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/S0315086013000396.

Boyd:2014:STE

[Boy14] John P. (John Philip) Boyd. Solving transcendental equations: the Chebyshev polynomial proxy and other numerical rootfinders, perturbation series, and oracles. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2014. ISBN 1-61197-351-1 (paperback). xviii + 460 pp. LCCN QA353.T7 B69 2014.

URL http://www.loc.gov/catdir/
enhancements/fy1503/2014017078-
b.html;   http://www.loc.gov/catdir/
enhancements/fy1503/2014017078-
d.html;   http://www.loc.gov/catdir/
enhancements/fy1503/2014017078-
t.html.

Banerjee:2011:LAM

[BR11] Sudipto Banerjee and Anindya Roy.

Linear Algebra and Matrix Analysis for Statistics. Chapman and Hall/CRC texts in statistical science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4200-9538-2 (hardback). xvii + 565 pp. LCCN QA184.2 .B36

2014. URL http://images.tandf.
co.uk/common/jackets/websmall/
978142009/9781420095388.jpg.

Brandt:2014:DAS

[Bra14] Siegmund Brandt. Data analysis: statistical and computational methods for scientists and engineers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2014. ISBN 3-319-03762-5 (ebook). ???? pp. LCCN QA273; QA273.

Baumgarte:2010:NRS

[BS10] Thomas W. Baumgarte and Stuart L.

(Stuart Louis) Shapiro. Numerical relativity: solving Einstein’s equations on the computer. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-51407-X. xviii + 698 pp. LCCN QC173.6 .B38 2010.

Borwein:2011:IMM

[BS11] Jonathan M. Borwein and Matthew P.

Skerritt. An introduction to modern mathematical computing. Springer Undergraduate Texts in Mathematics and Technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4614-0121-6. xvi + 216 pp. With MapleTM.

Brezinski:2013:WGV

[BS13] Claude Brezinski and Ahmed Sameh. Walter Gautschi, Volume 1: Selected Works with Commentaries. Contemporary Mathematicians. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-7033-1, 1-4614-7034-X. xi + 694 + 50 pp. LCCN

QA297. URL http://site.ebrary.
com/id/10787871.

Brezinski:2014:WGVa

[BS14a] Claude Brezinski and Ahmed Sameh, editors. Walter Gautschi. Volume 2: selected works with commentaries. Contemporary mathematicians. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2014. ISBN 1-4614-7048-X, 1-4614-7049-8 (e-book). xiii + 914 + 33 pp. LCCN QA404.5. URL http:/

/link.springer.com/10.1007/978-
1-4614-7049-6.

Brezinski:2014:WGVb

[BS14b] Claude Brezinski and Ahmed Sameh, editors. Walter Gautschi. Volume 3: selected works with commentaries. Contemporary mathematicians. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2014. ISBN 1-4614-7131-1, 1-4614-7132-X (e-book). xi + 767 + 91 + 29 pp. LCCN QA431. URL http:/

/link.springer.com/10.1007/978-
1-4614-7132-5.

Boyd:2018:IAL

[BV18] Stephen P. In. Boyd and Lieven Vandenberghe. Introduction to Applied Linear Algebra: Vectors, Matrices, and Least Squares. Cambridge University Press, Cambridge, UK, 2018. ISBN 1-108-69394-6, 1-316-51896-5 (hardcover). xii + 463 pp. LCCN QA184.2 .B69 2018.

Babuska:2011:FEI

[BWS11] Ivo Babuška, J. R. (John Robert)

Whiteman, and Theofanis Strouboulis. Finite elements: an introduction to the method and error estimation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-850670-8. xii + 323 pp. LCCN QA276.8 .B33X 2011.

URL http://www.loc.gov/catdir/
enhancements/fy1108/2010033235-
b.html;   http://www.loc.gov/catdir/
enhancements/fy1108/2010033235-
d.html;   http://www.loc.gov/catdir/
enhancements/fy1108/2010033235-

t.html. An introduction to the method and error estimation.

Chaitin-Chatelin:2012:EM

[CCAL12] Françoise Chaitin-Chatelin, Mario Ahués, and Walter Ledermann. Eigenvalues of matrices, volume 71 of Classics in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, revised edition, 2012. ISBN 1-61197-245-0. xxx + 410 pp. LCCN QA188 .C44 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1305/2012033049-
d.html;   http://www.loc.gov/catdir/
enhancements/fy1305/2012033049-
t.html;   http://www.loc.gov/catdir/
enhancements/fy1307/2012033049-
b.html.

Corless:2013:GIN

[CF13] Robert M. Corless and Nicolas Fillion. A Graduate Introduction to Numerical Methods: from the Viewpoint of Backward Error Analysis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-8452-9 (hardcover), 1-4614-8453-7 (e-book). xxxix + 868 pp. LCCN QA297 .C665 2013.

Chen:2011:FEM

[Che11] Zhangxin Chen. The finite element method: its fundamentals and applications in engineering. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 981-4350-56-7 (hardcover), 981-4350-57-5 (paperback). xxi + 326 pp. LCCN TA347.F5 C467 2011. Its fundamentals and applications in engineering.

Christensen:2010:FSE

[Chr10] Ole Christensen. Functions, spaces, and expansions: mathematical tools in physics and engineering. Applied and numerical harmonic analysis. Birkhäuser Boston Inc., Cambridge, MA, USA, 2010. ISBN 0-8176-4980-8. xix + 263 pp. LCCN QA331.7 .C57 2010.

Colonius:2014:DSL

[CK14] Fritz Colonius and Wolfgang Kliemann. Dynamical Systems and Linear Algebra, volume ume 158 of Graduate studies in mathematics. American Mathematical Society, Providence, RI, USA, 2014. ISBN 0-8218-8319-4. ???? pp. LCCN QA184.2 .C65 2014.

Cohen:2011:NAM

[Coh11] Harold Cohen. Numerical approximation methods. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-9836-5. xiv + 485 pp.

Datta:2010:NLA

[Dat10] Biswa Nath Datta. Numerical linear algebra and applications. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2010. ISBN 0-534-17466-3 (paperback), 0-89871-685-3. xxiv + 530 pp. LCCN QA184.2 .D38

2010.  URL http://www.gbv.de/
dms/ilmenau/toc/603672094.PDF;
 http://www.loc.gov/catdir/enhancements/
fy1001/2009025104-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1001/2009025104-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy1001/2009025104-t.html;   http:
//www.zentralblatt-math.org/zmath/
en/search/?an=1187.65027.

Davies:2011:FEM

[Dav11a] Alan J. Davies. The finite element method: an introduction with partial differential equations. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2011. ISBN 0-19-960913-6. ix + 297 pp. LCCN TA347.F5 D38 2011.

URL http://www.loc.gov/catdir/
enhancements/fy1211/2011022386-
b.html;   http://www.loc.gov/catdir/
enhancements/fy1211/2011022386-
d.html;   http://www.loc.gov/catdir/
enhancements/fy1211/2011022386-

t.html. An introduction with partial differential equations.

Davis:2011:MP

[Dav11b] Timothy A. Davis. MATLAB primer. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, eighth edition, 2011. ISBN 1-4398-2862-8. xvi + 232 pp. LCCN QA297 .D38 2011.

Davis:2012:LAP

[Dav12] Ernest Davis. Linear algebra and probability for computer science applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4665-0155-3 (hardcover). xviii + 413 pp. LCCN QA76.9.M35 D38 2012.

Dennis:2013:RSC

[Den13] Brian Dennis. The R student companion. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2013. ISBN 1-4398-7540-5 (paperback). xvii + 339 pp. LCCN QA276.45.R3 D46 2013.

                                        URL http://jacketsearch.tandf.
                                        co.uk/common/jackets/covers/websmall/
                                        978143987/9781439875407.jpg.

Dick:2010:DNS

[DP10] J. (Josef) Dick and Friedrich Pillichshammer. Digital nets and sequences: discrepancy and quasi-Monte Carlo integration. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19159-9 (hardback). xvii + 600 pp. LCCN QA298 .D53 2010. URL http:/

                                        /assets.cambridge.org/97805211/
                                        91593/cover/9780521191593.jpg.

Deuflhard:2011:NM

[DW11] Peter Deuflhard and Martin Weiser. Numerische Mathematik 3 (German) [Numerical mathematics 3]. de Gruyter Lehrbuch [de Gruyter Textbook]. Walter de Gruyter, New York, 2011. ISBN 3-11-021802-X. x + 432 pp. Adaptive Lösung partieller Differentialgleichungen. [Adaptive solutions of partial differential equations].

Eubank:2012:SCC

[EK12] Randall L. Eubank and Ana Kupresanin. Statistical computing in C++ and R. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4200-6650-1 (hardcover). xv + 540 pp. LCCN QA276.4 .E87 2012.

Etter:2010:IM

[Ett10] Delores M. Etter. Introduction to MATLAB. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2010. ISBN 0-13-608123-1. ???? pp. LCCN TA345 .E8724 2010.

Fiedler:2011:MGG

[Fie11] Miroslav Fiedler. Matrices and Graphs in Geometry, volume 139 of Encyclopedia of Mathematics and its Applications. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-46193-6. viii + 197 pp. LCCN QA447 .F45 2011. URL http:/

/assets.cambridge.org/97805214/
61931/cover/9780521461931.jpg;
http://catdir.loc.gov/catdir/
enhancements/fy1101/2010046601-
b.html;   http://catdir.loc.gov/
catdir/enhancements/fy1101/2010046601-
d.html;   http://catdir.loc.gov/
catdir/enhancements/fy1101/2010046601-
t.html.

Forster:2010:FSC

[FM10] Brigitte Forster and Peter Robert Massopust, editors. Four short courses on harmonic analysis: wavelets, frames, time-frequency methods, and applications to signal and image analysis. Applied and numerical harmonic analysis. Birkhäuser Boston Inc., Cambridge, MA, USA, 2010. ISBN 0-8176-4891-7, 0-8176-4890-9. xvii + 247 pp. LCCN QA403 .F68 2010. With contributions by Ole Christensen, Karlheinz Gröchenig, Demetrio Labate, Pierre Vandergheynst, Guido Weiss, and Yves Wiaux.

Fukuda:2010:MSI

[FVJT10] Komei Fukuda, Joris Van der Hoeven, Michael Joswig, and Nobuki Takayama, editors. Mathematical Software — ICMS 2010: Third International Congress on Mathematical Software, Kobe, Japan, September 13–17, 2010, Proceedings, Lecture Notes in Computer Science / Theoretical Computer Science and General Issues Ser. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-15581-2 (paperback). LCCN QA76.95 .I5654 2010.

URL http://link.springer.com/
openurl?genre=book&isbn=978-3-
642-15581-9.

Gautschi:2016:OPM

[Gau16] Walter Gautschi. Orthogonal polynomials in MATLAB: exercises and solutions, volume 26 of Software, environments, and tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2016. ISBN 1-61197-429-1 (paperback), 1-61197-430-5. ix + 335 pp. LCCN QA404.5 .G3564 2016.

Griffiths:2010:NMO

[GH10] David F. (David Francis) Griffiths and Desmond J. (Desmond J.) Higham.

Numerical methods for ordinary differential equations: initial value problems. Springer Undergraduate Mathematics Series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-85729-147-5. xiv + 271 pp. LCCN QA371 .G72 2010. Initial value problems.

Garcia:2017:SCL

[GH17] Stephan Ramon Garcia and Roger A.

Horn. A Second Course in Linear Algebra. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-10381-9 (hardcover). 442 (est.) pp. LCCN QA184.2 .G37 2017. URL http:/

/www.cambridge.org/us/academic/
subjects/mathematics/algebra/second-
course-linear-algebra?format=HB.

Graham:2012:NAM

[GHLS12] Ivan G. Graham, Thomas Y. Hou, Omar Lakkis, and Robert Scheichl, editors. Numerical Analysis of Multiscale Problems, volume 83 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSA6. ISBN 3-642-22060-6 (print), 3-642-22061-4 (e-book). ISSN 1439-7358. LCCN QA297 .N844 2012.

URL http://link.springer.com/
book/10.1007/978-3-642-22061-6;
 http://www.springerlink.com/content/
978-3-642-22061-6.  Ten invited

expository articles from the 91st LMS Durham Symposium on Numerical Analysis of Multiscale Problems, Durham, UK, 5–15 July 2010.

Golub:2010:MMQ

[GM10] Gene H. Golub and Gérard Meurant. Matrices, Moments and Quadrature with Applications. Princeton Series in Applied Mathematics. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-14341-2. xii + 363 pp.

Govil:2017:PAT

[GMQS17] Narendra Kumar Govil, Ram Mohapatra, Mohammed A. Qazi, and Gerhard Schmeisser, editors. Progress in Approximation Theory and Applicable Complex Analysis: In Memory of Q. I. Rahman, volume 117 of Springer Optimization and Its Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-49240-3 (print), 3-319-49242-X (e-book). LCCN QA402.5-402.6. URL https:

                                        //link.springer.com/chapter/10.
                                        1007/978-3-319-49242-1.

Gautschi:2011:ACH

[GMR11] Walter Gautschi, Giuseppe Mastroianni, and Themistocles M. Rassias, editors. Approximation and Computation: In Honor of Gradimir V. Milovanović, volume 42 of Springer Optimization and Its Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-6593-9 (paperback), 1-4419-6594-7 (e-book), 1-4419-6595-5, 1-4614-2703-7. xviii + 482 pp. LCCN QA39.2 .A67 2011; QA221 .A6345 2011.

Gilli:2011:NMO

[GMS11] Manfred Gilli, Dietmar Maringer, and Enrico Schumann, editors. Numerical Methods and Optimization in Finance. Elsevier Academic Press, Amsterdam, The Netherlands, 2011. ISBN 0-12-375662-6. xv + 584 pp. LCCN HG106 .G55 2011.

Green:2016:CMC

[Gre16a] Dan Green. Cosmology with MAT-LAB: with companion media pack. World Scientific Publishing Co. Pte. Ltd., Singapore, 2016. ISBN 981-310-839-8 (hardcover), 981-310-840-1 (paperback). xi + 250 pp. LCCN QB981 .G74 2016.

Greuel:2016:MSI

[Gre16b] Gert-Martin Greuel, editor. Mathematical Software — ICMS 2016: 5th International Conference, Berlin, Germany, July 11–14, 2016: proceedings, volume 9725 of Lecture Notes in Computer Science/Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-42431-9 (print), 3-319-42432-7 (electronic). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.9.M35. URL http://zbmath.
org/?q=an:1342.68017.

Gruber:2014:MAL

[Gru14] Marvin H. J. Gruber. Matrix Algebra for Linear Models. Wiley, New York, NY, USA, 2014. ISBN 1-118-59255-7 (hardcover), 1-118-60881-X (e-book), 1-118-60874-7 (e-book), 1-118-80041-9 (e-book). xv + 375 pp. LCCN QA279 .G78 2014.

Griffiths:2012:TWA

[GS12] Graham W. Griffiths and W. E.

Schiesser. Traveling wave analysis of partial differential equations: numerical and analytical methods with MAT-LAB and Maple. Academic Press, New York, NY, USA, 2012. ISBN 0-12-384652-8 (hardcover). xiii + 447 pp. LCCN QA374 .G75 2012.

Graham:2013:SSM

[GT13] C. (Carl) Graham and D. (Denis) Talay. Stochastic Simulation and Monte Carlo Methods: Mathematical Foundations of Stochastic Simulation, volume 68 of Stochastic Modelling and Applied Probability. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-39362-4. ISSN 0172-4568. xvi + 260 + 4 pp. LCCN QA273.A1-274.9; QA274-274.9.

Gustafsson:2011:FSC

[Gus11] Bertil Gustafsson. Fundamentals of scientific computing, volume 8 of Texts in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-19494-X. xiv + 316 pp.

Galvis:2011:IAN

[GV11] Juan Galvis and Henrique Versieux. Introdução à aproximação numérica de equações diferenciais parciais via o método de elementos finitos. (Portuguese) [Introduction to numerical approximation of partial differential equations via the finite-element method]. Publicações Matemáticas do IMPA. [IMPA Mathematical Publications]. Instituto Nacional de Matemática Pura e Aplicada (IMPA), Rio de Janeiro, Brasil, 2011. ISBN 85-244-0325-X. 91 pp. 28o Colóquio Brasileiro de Matemática. [28th Brazilian Mathematics Colloquium].

Golub:2013:MC

[GV13] Gene H. Golub and Charles F. Van Loan. Matrix Computations. Johns Hopkins Studies in the Mathematical Sciences. The Johns Hopkins University Press, Baltimore, MD, USA, fourth edition, 2013. ISBN 1-4214-0794-9 (hardcover), 1-4214-0859-7 (ebook). xxi + 756 pp. LCCN QA188 .G65 2013. URL https:/

/jhupbooks.press.jhu.edu/title/
matrix-computations.

Gander:2012:ERG

[GW12] Martin J. Gander and Gerhard Wanner. From Euler, Ritz, and Galerkin to modern computing. SIAM Review, 54(4):627–666, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hamming:2012:IAN

[Ham12] R. W. (Richard Wesley) Hamming. Introduction to Applied Numerical Analysis. Dover books on mathematics. Dover, New York, NY, USA, 2012. ISBN 0-486-48590-0 (paperback). x + 331 pp. LCCN QA297 .H275 2012.

Hassanieh:2018:SFT

[Has18] Haitham Hassanieh. The Sparse Fourier Transform: Theory and Practice, volume 19 of ACM book series. ACM Press, New York, NY 10036, USA, 2018. ISBN 1-947487-07-8 (hardcover), 1-947487-04-3 (paperback), 1-947487-05-1 (e-book). xvii + 260 pp. LCCN QC20.7.F67 H37 2018.

Hermann:2011:NM

[Her11] Martin Hermann. Numerische Mathematik. R. Oldenbourg, München, Germany, expanded edition, 2011. ISBN 3-486-70820-1. xiv + 563 pp.

Hollig:2013:AMB

[HH13] Klaus Höllig and Jörg Hörner. Approximation and modeling with Bsplines. Applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2013. ISBN 1-61197-294-9. xiii + 214 pp. LCCN QA224 .H645 2013.

Hanson:2014:NCM

[HH14] Richard J. Hanson and Tim Hopkins. Numerical computing with modern Fortran. Applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2014. ISBN 1-61197-311-2 (paperback), 1-61197-312-0 (e-book). xv + 244 pp. LCCN QA76.73.F25 H367 2013.

Higham:2017:MG

[HH17] Desmond J. Higham and Nicholas J. Higham. MATLAB guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2017. ISBN 1-61197-465-8. xxvi + 476 pp.

Hill:2015:LSP

[Hil15] Christian Hill. Learning scientific programming with Python. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-07541-6 (hardcover), 1-107-42822-X (paperback). vii + 452 pp. LCCN Q183.9 .H58 2015.

Horn:2012:MA

[HJ12] Roger A. Horn and Charles R. (Charles Royal) Johnson. Matrix Analysis. Cambridge University Press, Cambridge, UK, second edition, 2012. ISBN 0-521-83940-8 (hardcover), 0-521-54823-3 (paperback), 1-283-74139-3, 1-139-77904-4, 1-139-77600-2 (ebook), 1-139-02041-2 (e-book). xviii + 643 pp. LCCN QA188 .H66 2012.

Hunt:2014:GMB

[HLR14] Brian R. Hunt, Ronald L. Lipsman, and Jonathan M. (Jonathan Micah) Rosenberg. A guide to MATLAB: for beginners and experienced users: updated for MATLAB 8 and Simulink 8. Cambridge University Press, Cambridge, UK, third edition, 2014. ISBN 1-107-66222-2 (paperback). ???? pp. LCCN QA297 .H86 2014.

Hogben:2014:HLA

[Hog14] Leslie Hogben, editor. Handbook of Linear Algebra. Discrete mathematics and its applications. CRC Press/Taylor and Francis Group, Boca Raton, FL, USA, second edition, 2014. ISBN 1-4665-0728-4 (hardcover). ???? pp. LCCN QA184.2 .H36 2014.

URL http://marc.crcnetbase.com/
isbn/9781466507296.

Howard:2017:CMN

[How17] James Patrick Howard, II. Computational methods, for numerical analysis with R. CRC Press/Taylor and Francis Group, Boca Raton, FL, USA, 2017. ISBN 1-4987-2363-2 (hardcover), 1-4987-2364-0 (ebook), 1-315-12019-4 (e-book). xx + 257 pp. LCCN QA297 .H67

2017. URL http://www.crcnetbase.
com/isbn/9781498723640.

Hansen:2013:LSD

[HPS13] Per Christian Hansen, V. (Víctor) Pereyra, and Godela Scherer. Least Squares Data Fitting with Applications. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0786-8 (hardcover), 1-4214-0858-9 (e-book). xv + 305 pp. LCCN QA275 .H26

                                        2013. URL http://muse.jhu.edu/
                                        books/9781421408583/.

Hilber:2013:CMQ

[HRSW13] Norbert Hilber, Oleg Reichmann, Ch. (Christoph) Schwab, and Christoph Winter. Computational Methods for Quantitative Finance: Finite Element Methods for Derivative Pricing. Springer Finance. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-35400-9. ISSN 1616-0533. xiii + 299 + 57 pp. LCCN QA273.A1-274.9; QA274-274.9.

Jones:2014:ISP

[JMR14] Owen (Owen Dafydd) Jones, Robert Maillardet, and Andrew (Andrew P.) Robinson. Introduction to scientific programming and simulation using R. The R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2014. ISBN 1-4665-6999-9, 1-4665-7001-6. xxiv + 582 pp. LCCN Q183.9 .J65 2014.

Johnson:2011:EMS

[Joh11] Richard K. Johnson. The elements of MATLAB style. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-73258-1. ???? pp. LCCN QA76.73.M296 J64 2011.

Klee:2011:SDS

[KA11] Harold Klee and Randal Allen. Simulation of dynamic systems with MAT-LAB(R) and Simulink(R). CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2011. ISBN 1-4398-3673-6 (hardback). xix + 795 pp. LCCN QA76.9.C65 K585 2011. With a foreword by Chris Bauer and Chris Schwarz.

Kurzak:2010:SCM

[KBD10] Jakub Kurzak, David A. Bader, and J. J. Dongarra, editors. Scientific computing with multicore and accelerators, volume 10 of Chapman and Hall/CRC computational science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-2536-X (hardback). xxxiii + 480 pp. LCCN Q183.9 .S325 2010.

Kushner:2014:NMS

[KD14] Harold J. Kushner and Paul Dupuis.

Numerical Methods for Stochastic Control Problems in Continuous Time, volume 24 of Stochastic Modelling and Applied Probability. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2014. ISBN 1-4612-6531-2. ISSN 0172-4568. xii + 476 pp. LCCN QA273.A1-274.9; QA274-274.9.

Kepner:2011:GAL

[KG11] Jeremy V. Kepner and J. R. (John R.)

Gilbert. Graph algorithms in the language of linear algebra. Software, environments, and tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2011. ISBN 0-89871-990-9 (hardcover). xxvii + 361 pp. LCCN QA166.245 .K47 2011.

                                        URL http://www.loc.gov/catdir/
                                        enhancements/fy1113/2011003774-
                                        b.html;   http://www.loc.gov/catdir/
                                        enhancements/fy1113/2011003774-
                                        d.html;   http://www.loc.gov/catdir/
                                        enhancements/fy1113/2011003774-
                                        t.html.

Kharab:2012:INM

[KG12] Abdelwahab Kharab and Ronald B. Guenther. An introduction to numerical methods: a MATLAB approach. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2012. ISBN 1-4398-6899-9 (hardback), 1-4398-6900-6 (e-book). 14 + 567 pp. LCCN QA297 .K52 2012.

Kiusalaas:2010:NMEa

[Kiu10a] Jaan Kiusalaas. Numerical methods in engineering with MATLAB(R). Cambridge University Press, Cambridge, UK, second edition, 2010. ISBN 0-521-19133-5 (hardback). xi + 431 pp. LCCN TA345 .K58 2010.

Kiusalaas:2010:NMEb

[Kiu10b] Jaan Kiusalaas. Numerical methods in engineering with Python. Cambridge University Press, Cambridge, UK, second edition, 2010. ISBN 0-521-19132-7 (hardcover), 0-511-67694-8 (ebook). x + 422 pp. LCCN TA345 .K584 2010.

Kiusalaas:2013:NME

[Kiu13] Jaan Kiusalaas. Numerical methods in engineering with Python 3. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-03385-3. xi + 423 pp. LCCN TA345 .K584 2013.

King:2010:NSM

[KM10a] Michael R. King and Nipa A. Mody.

Numerical and statistical methods for bioengineering: applications in MAT-LAB. Cambridge texts in biomedical engineering. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-87158-1, 0-511-90984-5 (e-book). xii + 581 pp. LCCN R857.M34 K56 2010eb. URL http:/

/assets.cambridge.org/97805218/
71587/cover/9780521871587.jpg;
 http://site.ebrary.com//lib/yale/
Doc?id=10431397.

Kornerup:2010:FPN

[KM10b] Peter Kornerup and David W.

Matula. Finite Precision Number Systems and Arithmetic, volume 133 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-76135-2 (hardcover). xv + 699 pp. LCCN QA248 .K627 2010. URL http:/

/assets.cambridge.org/97805217/
61352/cover/9780521761352.jpg;
 http://catdir.loc.gov/catdir/
enhancements/fy1011/2010030521-
b.html;   http://catdir.loc.gov/
catdir/enhancements/fy1011/2010030521-
d.html;   http://catdir.loc.gov/
catdir/enhancements/fy1011/2010030521-
t.html.

King:2011:NSM

[KM11] Michael R. King and Nipa A. Mody.

Numerical and statistical methods for bioengineering: applications in MAT-LAB. Cambridge Texts in Biomedical Engineering. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-87158-1 (hardback). xii + 581 pp.

LCCN R857.M34 K56 2011; R857.M34 K56X 2011 (LC). Applications in MATLAB.

Kneusel:2016:NC

[Kne16] Ronald T. Kneusel. Numbers and Computers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-35940-1 (softcover), 3-319-17260-3 (ebook). xi + 231 pp. LCCN QA241 .K54 2016.

Kneusel:2017:NC

[Kne17] Ronald T. Kneusel. Numbers and Computers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2017. ISBN 3-319-50507-6, 3-319-50508-4 (e-book). xiii + 346 pp. LCCN

                                        ???? URL http://link.springer.
                                        com/10.1007/978-3-319-50508-4.

Kilmer:2010:GWS

[KO10] Misha Elena Kilmer and Dianne P. O’Leary, editors. G. W. Stewart: selected works with commentaries. Contemporary mathematicians. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2010. ISBN 0-8176-4967-0 (hardcover), 0-8176-4968-9 (e-book). xii + 729 pp. LCCN QA188 .S74 2010.

                                        URL  http://public.eblib.com/
                                        EBLPublic/PublicView.do?ptiID=
                                        64586;   http://rave.ohiolink.
                                        edu/ebooks/ebc/978081764968;
                                        http://site.ebrary.com/id/1042123.

Kutz:2013:DDM

[Kut13] Jose Nathan Kutz. Data-driven modeling and scientific computation: methods for complex systems and big data. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966033-6 (hardcover), 0-19-966034-4 (paperback). xvii + 638 pp. LCCN Q183.9 .K88 2013.

Kugler:2012:AMB

[KW12] Philipp Kügler and Wolfgang Windsteiger. Algorithmische Methoden. Band 2. Mathematik Kompakt. [Compact Mathematics]. Birkhäuser/Springer Basel AG, Basel, 2012. ISBN 3-7643-8515-4; 3-7643-8516-2. viii + 159 pp. Funktionen, Matrizen, multivariate Polynome. [Functions, matrices, multivariate polynomials].

Lange:2010:NAS

[Lan10] Kenneth Lange. Numerical analysis for statisticians. Statistics and Computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2010. ISBN 1-4419-5944-0 (hardcover). xvi + 604 pp. LCCN QA297 .L34 2010.

Langtangen:2012:PSP

[Lan12] Hans Petter Langtangen. A primer on scientific programming with Python, volume 6 of Texts in computational science and engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2012. ISBN 3-642-30292-0, 3-642-30293-9 (e-book). ISSN 1611-0994. LCCN QA76.73.P98

L36 2012. URL http://site.ebrary.
com/id/10650410.

Laub:2012:CMA

[Lau12] Alan J. Laub. Computational matrix analysis. Other titles in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2012. ISBN 1-61197-220-5 (paperback), 1-61197-221-3 (e-book). xiii + 154 pp. LCCN QA274.2 .L38 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1211/2011050702-
b.html;   http://www.loc.gov/catdir/
enhancements/fy1211/2011050702-
d.html;   http://www.loc.gov/catdir/
enhancements/fy1211/2011050702-
t.html.

Lay:2014:LAAb

[Lay14] David C. Lay. Linear Algebra and its Applications. Pearson Education Limited, Harlow, Essex, fourth edition, 2014. ISBN 1-292-02055-5. ii + 784 pp. LCCN ????

Larson:2013:FEM

[LB13] Mats G. Larson and Fredrik Bengzon. The finite element method: theory, implementation, and applications, volume 10 of Texts in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-33286-2, 3-642-33287-0 (ebook). ISSN 1611-0994. xviii + 385 pp. LCCN TA347.F5 L37 2013.

Li:2018:NSD

[LQT18] Zhilin Li, Zhonghua Qiao, and Tao Tang. Numerical solution of differential equations: introduction to finite difference and finite element methods. Cambridge University Press, Cambridge, UK, 2018. ISBN 1-107-16322-6 (hardcover), 1-316-61510-3 (paperback), 1-316-67872-5. ix + 293 pp. LCCN QA371 .L59 2018.

Layton:2012:ADM

[LR12] William J. Layton and Leo G. Rebholz. Approximate Deconvolution Models of Turbulence: Analysis, Phenomenology and Numerical Analysis, volume 2042 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LN-MAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 184 pp. LCCN QA3 .L28 no. 2042.

URL http://link.springer.com/
book/10.1007/978-3-642-24409-4;
 http://www.springerlink.com/content/
978-3-642-24409-4.

Magoules:2010:FGC

[Mag10] F. (Frédéric) Magoulès, editor. Fundamentals of grid computing: theory, algorithms and technologies. Chapman and Hall/CRC numerical analysis and scientific computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-0367-6 (hardcover). xxi + 298 pp. LCCN QA76.9.C58 F86 2010.

Mehta:2015:MPS

[Meh15] Hemant Kumar Mehta. Mastering Python scientific computing: a complete guide for Python programmers to master scientific computing using Python APIs and tools. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78328-883-3, 1-78328-882-5. ???? pp. LCCN QA76.73.P98. URL http://

proquest.safaribooksonline.com/
?fpi=9781783288823.

Miller:2014:NAE

[Mil14] G. Miller. Numerical Analysis for Engineers and Scientists. Cambridge University Press, Cambridge, UK, 2014.

ISBN 1-107-02108-1 (hardcover), 1-139-10818-2 (ebook). x + 572 pp. LCCN QA297 .M55 2014.

Meckes:2018:LA

[MM18] Elizabeth S. Meckes and Mark W. Meckes. Linear Algebra. Cambridge mathematical textbooks. Cambridge University Press, Cambridge, UK, 2018. ISBN 1-107-17790-1 (hardcover). xvi + 427 pp. LCCN QA184.2 .M43 2018.

Martinez:2017:EDA

[MMS17] Wendy L. Martinez, Angel R. Martinez, and Jeffrey L. Solka. Exploratory data analysis with MATLAB, volume 4 of Computer science and data analysis series. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2017. ISBN 1-4987-7606-X (hardcover), 1-315-33081-4 (Mobi e-book), 1-4987-7607-8 (PDF e-book), 1-315-34984-1 (ePub). xv + 590 pp. LCCN QA278 .M3735 2017.

Moin:2010:FEN

[Moi10] Parviz Moin. Fundamentals of engineering numerical analysis. Cambridge University Press, Cambridge, UK, second edition, 2010. ISBN 0-521-88432-2 (hardcover), 0-521-71123-1. xiv + 241 pp. LCCN TA335 .M65 2010.

Molitierno:2012:ACM

[Mol12] Jason J. Molitierno. Applications of combinatorial matrix theory to Laplacian matrices of graphs. Discrete mathematics and its applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-6337-7 (hardcover). 405 pp. LCCN QA166.243 .M65 2012.

Monahan:2011:NMS

[Mon11] John F. Monahan. Numerical methods of statistics. Cambridge Series in Statistical and Probabilistic Mathematics. Cambridge University Press, Cambridge, UK, second edition, 2011. ISBN 0-521-13951-1 (paperback), 0-521-19158-0. xvi + 447 pp. LCCN QA276.4 .M65 2011 (LC); QA276.4 .M65 2011.

Muller:2016:EFA

[Mul16] Jean-Michel Muller. Elementary Functions: Algorithms and Implementation. Birkhäuser Boston Inc., Cambridge, MA, USA, third edition, 2016. ISBN 1-4899-7981-6 (print), 1-4899-7983-2 (e-book). xxv + 283 pp. LCCN QA331 .M866 2016.

Nahin:2011:NCT

[Nah11] Paul J. Nahin. Number-crunching:

taming unruly computational problems from mathematical physics to science fiction. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14425-7 (hardcover), 1-4008-3958-0 (e-book). xxvi + 376 pp. LCCN QC20.7.E4 N34

2011. URL http://www.jstor.org/
stable/10.2307/j.ctt7rk7v.

Naumann:2011:ADC

[Nau11] Uwe Naumann. The art of differentiating computer programs: an introduction to algorithmic differentiation. Software, environments, and tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2011. ISBN 1-61197-206-X. xviii + 340 pp. LCCN QA76.76.A98 N38 2011.

URL http://www.loc.gov/catdir/
enhancements/fy1201/2011032262-

                                        b.html;   http://www.loc.gov/catdir/
                                        enhancements/fy1201/2011032262-
                                        d.html;   http://www.loc.gov/catdir/
                                        enhancements/fy1201/2011032262-
                                        t.html.

Nassif:2014:INA

[NF14] Nabil Nassif and Dolly Khuwayri Fayyad. Introduction to Numerical Analysis and Scientific Computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-8948-5 (hardcover). xix + 311 pp. LCCN QA297 .N37 2014.

Nakao:2019:NVM

[NPW19] Mitsuhiro T. Nakao, Michael Plum, and Yoshitaka Watanabe. Numerical Verification Methods and Computer-assisted Proofs for Partial Differential Equations, volume 53 of Springer series in computational mathematics. Springer, Singapore, 2019. ISBN 981-13-7668-9, 981-13-7669-7 (e-book). ISSN 0179-3632. xiii + 467 pp. LCCN QA377.

                                        URL https://www.springer.com/
                                        gp/book/9789811376689.

Ochsner:2013:ODF

[ÖM13] Andreas Öchsner and Markus Merkel. One-dimensional finite elements: an introduction to the FE method. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-31796-0 (hardcover); 3-642-31797-9 (e-book). xxiii + 398 pp. LCCN TA347.F5 O24 2013. An introduction to the FE method.

Onate:2010:SAF

[Oña10] Eugenio Oñate. Structural analysis with the finite element method: linear statistics. Lecture notes on numerical methods in engineering and sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-4020-8733-0. xxiv + 472 pp. LCCN TA347.F5 O63 2009.

Oberkampf:2010:VVS

[OR10] William L. Oberkampf and Christopher J. Roy. Verification and validation in scientific computing. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-11360-1. ???? pp. LCCN Q183.9 .O24 2010. URL http:/

/assets.cambridge.org/97805211/
13601/cover/9780521113601.jpg.

Pozrikidis:2013:XSC

[Poz13] C. Pozrikidis. XML in scientific computing. Chapman and Hall/CRC numerical analysis and scientific computing series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4665-1227-X (hardback). xv + 243 pages pp. LCCN Q183.9 .P69 2013.

Pont:2012:DDW

[PR12] Jean-Claude Pont and Christophe Rossel. Le destin douloureux de Walther Ritz (1878–1909), physicien théoricien de génie, volume 24 of Cahiers de Vallesia. Archives de l’état du Valais, Vallesia, France, 2012. ISBN 2-9700636-5-4 (hardcover). 264 + 41 pp. LCCN ????

Quarteroni:2010:SCM

[QSG10] Alfio Quarteroni, Fausto Saleri, and Paola Gervasio. Scientific computing with Matlab and Octave, volume 2 of Texts in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-12429-1. xvi + 360 pp. LCCN ???? Third edition [of MR2253397].

Quarteroni:2014:SCM

[QSG14] Alfio Quarteroni, Fausto Saleri, and Paola Gervasio. Scientific computing with Matlab and Octave, volume 2 of Texts in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-642-45366-X (hard cover). xviii + 450 (est.) pp. LCCN ???? URL http://

                                        link.springer.com/book/10.1007/
                                        978-3-642-45367-0.

Razavy:2011:HQM

[Raz11] Mohsen Razavy. Heisenberg’s quantum mechanics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 981-4304-11-5 (paperback), 981-4304-10-7. xix + 657 pp. LCCN QC174.12 .R39 2011.

Robert:2010:IMC

[RC10] Christian P. Robert and George Casella. Introducing Monte Carlo methods with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-4419-1575-3 (paperback), 1-4419-1576-1 (ebk.). xix + 283 pp. LCCN QA298 .R63 2010.

Rebaza:2012:FCA

[Reb12] Jorge Rebaza. A first course in applied mathematics. Wiley, New York, NY, USA, 2012. ISBN 1-118-22962-2. xvi + 439 pp. LCCN TA342 .R43 2012.

URL http://www.loc.gov/catdir/
enhancements/fy1201/2011043340-
d.html;   http://www.loc.gov/catdir/
enhancements/fy1201/2011043340-
t.html;   http://www.loc.gov/catdir/
enhancements/fy1210/2011043340-
b.html.

Rizopoulos:2012:JML

[Riz12] Dimitris Rizopoulos. Joint models for longitudinal and time-to-event data: with applications in R, volume 6 of Chapman and Hall/CRC biostatistics series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-7286-4 (hardcover). xiv + 261 pp. LCCN QA279 .R59 2012.

Rossant:2013:LII

[Ros13] Cyrille Rossant. Learning IPython for interactive computing and data visualization: Learn IPython for interactive Python programming, highperformance numerical computing, and data visualization. Open source: community experience distilled. Packt Publishing, Birmingham, UK, 2013. ISBN 1-78216-993-8 (paperback), 1-78216-994-6 (e-book), 1-299-54508-4 (e-book). iv + 123 pp. LCCN QA76.73.P98 .R677 2013.

Rossant:2014:IIC

[Ros14] Cyrille Rossant. IPython interactive computing and visualization cookbook: over 100 hands-on recipes to sharpen your skills in high-performance numerical computing and data science with Python. Packt Publishing Ltd., Birmingham, UK, 2014. ISBN 1-78328-481-1, 1-78328-482-X (e-book), 1-322-16622-6 (e-book). v + 494 pp. LCCN QA76.73.P98 R677 2014.

Reich:2015:IPO

[RZ15] Simeon Reich and Alexander J. Zaslavski, editors. Infinite products of operators and their applications: a research workshop of the Israel Science Foundation: May 21–24, 2012, Haifa, Israel: Israel mathematical conference proceedings, volume 636 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 2015. ISBN 1-4704-1480-5 (paperback). LCCN QA329 .I54

                                        2015. URL http://www.ams.org/
                                        books/conm/636/.

Romisch:2016:MAM

[RZ16] Werner Römisch and Thomas Zeugmann. Mathematical Analysis and the Mathematics of Computation. Springer, Cham, Switzerland, 2016. ISBN 3-319-42753-9 (hardcover), 3-319-42755-5 (e-book). xxiii + 704

                                        pp. URL http://link.springer.
                                        com/10.1007/978-3-319-42755-3.

Saad:2011:NML

[Saa11] Youcef Saad. Numerical Methods for Large Eigenvalue Problems, volume 66 of Classics in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2011. ISBN 1-61197-072-5. xv + 276 pp. LCCN QA188 .S18

                                        2011. URL http://www.cs.umn.edu/
                                        ~saad/eig_book_2ndEd.pdf.

Scherzer:2015:HMM

[Sch15] Otmar Scherzer, editor. Handbook of Mathematical Methods in Imaging. SpringerReference, New York, NY, USA, second edition, 2015. ISBN 1-4939-0789-1 (set), 1-4939-0790-5 (ebook). xviii + 2178 (3 volumes) pp. LCCN RC78.7.D53 H358 2015.

Schiesser:2017:SCM

[Sch17] William E. Schiesser. Spline Collocation Methods for Partial Differential Equations: with Applications in R. Wiley, New York, NY, USA, 2017. ISBN 1-119-30103-3 (hardcover), 1-119-30105-X (PDF), 1-119-30104-1 (ePub), 1-119-30106-8 (online). xv + 549 pp. LCCN QA377 .S355 2017.

Soetaert:2012:SDE

[SCM12] Karline Soetaert, Jeff Cash, and Francesca Mazzia. Solving differential equations in R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-28069-2 (hardcover), 3-642-28070-6 (e-book). xvi + 248 pp. LCCN QA371.5.D37 S64 2012.

Scott:2011:NA

[Sco11] L. Ridgway Scott. Numerical analysis.

Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14686-1 (hardcover). xiv + 325 pp. LCCN QA297 .S38 2011; QA297 .S393 2011.

Stakgold:2011:GFB

[SH11] Ivar Stakgold and Michael J. Holst.

Green’s Functions and Boundary Value Problems, volume 99 of Pure and applied mathematics. Wiley, New York, NY, USA, third edition, 2011. ISBN 0-470-60970-2 (hardcover). xxi + 855 pp. LCCN QA379 .S72 2011.

Sirca:2012:CMP

[ŠH12] Simon Širca and Martin Horvat.

Computational methods for physicists.

Graduate Texts in Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-32477-0; 3-642-32478-9. xx + 715 pp. Compendium for students.

Shapira:2012:SPC

[Sha12] Yair Shapira. Solving PDEs in C++: numerical methods in a unified objectoriented approach. Computational science and engineering series. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2012. ISBN 1-61197-216-7 (paperback). xxxii + 776 pp. LCCN QA377 .S466 2012.

Singh:2013:LAS

[Sin13] Kuldeep Singh. Linear Algebra: Step by Step. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-965444-1 (paperback), 0-19-150776-8 (e-book). viii + 608 pp. LCCN QA184.2 .S56 2014.

Stys:2014:LNN

[SS14] Tadeusz Styś and Krystyna Styś. Lecture Notes in Numerical Analysis with Mathematica. Bentham Science Publishers, Inc., Sharjah, United Arab Emirates, 2014. ISBN 1-60805-942-1 (e-book), 1-60805-943-X. 243 pp. LCCN QA298 .S797 2014. URL http://

ebooks.benthamsciencepublisher.
org/book/9781608059423/.

SouzadeCursi:2015:UQS

[SS15] Eduardo Souza de Cursi and Rubens Sampaio. Uncertainty Quantification and Stochastic Modeling with Matlab. ISTE Press Ltd, London, UK, 2015. ISBN 0-08-100471-0 (ebook), 1-78548-005-7. LCCN QA274.2.

URL  http://alltitles.ebrary.
com/Doc?id=11040161;   http://
lib.myilibrary.com?id=762946;
http://public.eblib.com/choice/
PublicFullRecord.aspx?p=2007484;
 http://www.sciencedirect.com/
science/book/9781785480058.

Salehi:2018:NISa

[SS18a] Younes Salehi and William E.

Schiesser. Numerical integration of space fractional partial differential equations. Volume 1, Introduction to algorithms and computer coding in R, volume 19 of Synthesis lectures on mathematics and statistics. Morgan and Claypool Publishers, San Rafael, CA, USA, 2018. ISBN 1-68173-207-6 (paperback), 1-68173-208-4 (e-book). ISSN 1938-1743 (print), 1938-1751 (electronic). xii + 189 pp. LCCN QA372 .S266 2018.

Salehi:2018:NISb

[SS18b] Younes Salehi and William E.

Schiesser. Numerical integration of space fractional partial differential equations. Volume 2, Applications from classical integer PDEs, volume 20 of Synthesis lectures on mathematics and statistics. Morgan and Claypool Publishers, San Rafael, CA, USA, 2018. ISBN 1-68173-209-2 (hardcover), 1-68173-210-6 (e-book). ISSN 1938-1743 (print), 1938-1751 (electronic). xii + 183–375 pp. LCCN QA372 .S2662 2018.

Stenger:2011:HSN

[Ste11] Frank Stenger. Handbook of Sinc Numerical Methods. Chapman and Hall/CRC numerical analysis and scientific computation series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-2158-5 (hardback), 1-4398-2159-3 (e-book). xx + 463 pp. LCCN QA372 .S8195 2010. URL http:/

                                        /www.crcpress.com/product/isbn/
                                        9781439821589.

Stewart:2014:PS

[Ste14] John Stewart. Python for scientists. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-06139-3 (hardcover), 1-107-68642-3. ???? pp. LCCN Q183.9 .S865 2014. URL http:/

                                        /assets.cambridge.org/97811070/
                                        61392/cover/9781107061392.jpg.

Snieder:2015:GTM

[SV15] Roel Snieder and Kasper Van Wijk. A Guided Tour of Mathematical Methods for the Physical Sciences. Cambridge University Press, Cambridge, UK, third edition, 2015. ISBN 1-107-08496-2 (hardcover), 1-107-64160-8 (paperback), 1-139-01354-8 (e-book). xxii + 560 pp. LCCN QA300 .S794 2015.

Temme:2015:AMI

[Tem15] Nico M. Temme. Asymptotic Methods for Integrals, volume 6 of Series in Analysis. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2015. ISBN 981-4612-15-4 (hardcover), 981-4612-16-2 (e-book). xxii + 605 pp.

Tveito:2010:ESC

[TLNC10] Aslak Tveito, Hans Petter Langtangen, Bjørn Frederik Nielsen, and Xing Cai. Elements of scientific computing: with 88 figures and 18 tables, volume 7 of Texts in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-11298-6. ISSN 1611-0994. xii + 459 pp. LCCN Q183.9 .E446 2010.

Trappenberg:2010:FCN

[Tra10] Thomas P. Trappenberg. Fundamentals of computational neuroscience. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2010. ISBN 0-19-956841-3 (paperback). xxv + 390 pp. LCCN QP357.5 .T746 2010.

Trefethen:2013:ATA

[Tre13] Lloyd N. Trefethen. Approximation Theory and Approximation Practice. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2013. ISBN 1-61197-239-6 (paperback). viii + 305 pp. LCCN QA221 .T73 2013.

Tucker:2011:VNS

[Tuc11] Warwick Tucker. Validated numerics:

a short introduction to rigorous computations. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14781-7 (hardcover). ???? pp. LCCN QA76.95 .T83 2011.

VanLoan:2010:ITC

[VF10] Charles F. Van Loan and K.-Y. Daisy Fan. Insight through computing: a MATLAB introduction to computational science and engineering. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2010. ISBN 0-89871-691-8. xviii + 434 pp. LCCN QA297 .V25 2010.

                                        URL http://www.loc.gov/catdir/
                                        enhancements/fy1007/2009030277-
                                        b.html;   http://www.loc.gov/catdir/
                                        enhancements/fy1007/2009030277-
                                        d.html;   http://www.loc.gov/catdir/
                                        enhancements/fy1007/2009030277-
                                        t.html.

Anonymous:2014:NMO

[Wal14] Eric Walter. Numerical Methods and Optimization: a Consumer Guide. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-07670-1, 3-319-07671-X (e-book). xv + 476 pp. LCCN QA402.5 .W358 2014eb.

                                        URL http://www.loc.gov/catdir/
                                        enhancements/fy1411/2014940746-
                                        d.html;   http://www.loc.gov/catdir/
                                        enhancements/fy1411/2014940746-
                                        t.html.

Wartak:2013:CPI

[War13] Marek S. Wartak. Computational photonics: an introduction with MATLAB. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-00552-3 (hardcover). xiii + 452 pp. LCCN TK8304 .W37 2013. URL http:/

                                        /assets.cambridge.org/97811070/
                                        05525/cover/9781107005525.jpg.

Watkins:2010:FMC

[Wat10] David S. Watkins. Fundamentals of Matrix Computations. Pure and applied mathematics. Wiley, New York, NY, USA, third edition, 2010. ISBN 0-470-52833-8 (hardcover). xvi + 644 pp. LCCN QA188 .W38 2010.

Watkins:2011:FA

[Wat11] David S. Watkins. Francis’s algorithm. American Mathematical Monthly, 118(5):387–403, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.
org/stable/info/10.4169/amer.math.
monthly.118.05.387.

Wulf:2012:CVR

[Wul12] Andrea Wulf. Chasing Venus: the race to measure the heavens. Alfred A. Knopf, New York, NY, USA, 2012. ISBN 0-307-70017-8 (hardcover), 0-307-95861-2 (e-book). xxvi + 304 pp. LCCN QB205.A2 W85 2012.

Zhang:2011:MTB

[Zha11] Fuzhen Zhang. Matrix Theory: Basic Results and Techniques. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2011. ISBN 1-4614-1098-3 (paperback), 1-4614-1099-1 (e-book). xviii + 399 pp. LCCN QA188 .Z47 2011.

URL http://www.loc.gov/catdir/
enhancements/fy1406/2011935372-
b.html;   http://www.loc.gov/catdir/
enhancements/fy1406/2011935372-
d.html;   http://www.loc.gov/catdir/
enhancements/fy1406/2011935372-
t.html.